OpenStudy (anonymous):

2/3 times 5 and 1/2.

OpenStudy (anonymous):

2/3 * 5 *1/2 = 10/3 * 1/2= 10/6

OpenStudy (anonymous):

Step 1 : Convert all numbers to fractions. Step 2 : 5 and one half is the fraction 11/2, (called an improper fraction). Step 3 : Multiply the tops of the two fractions together, ie, 2 x 11 = 22. Step 4 : Multiply the bottoms of the two fractions together, ie, 3 x 2 = 6. Step 5 : Answer is 22/6 but you must make this fraction look smaller so, divide the top and bottom by 2. Step 6 : Answer is 11/3 or, if you want to put it in whole number plus fraction form, 3 and 2/3.